OpenStudy (anonymous):

one more for ya if your up for it, need all solutions in the form of (solution+2Kpi) Find all solutions of the equation: cos^2(x)+5sin(x)+5=0 Thank you in advance

OpenStudy (anonymous):

what's this?

OpenStudy (anonymous):

algorithm for basic trig equations

OpenStudy (anonymous):

for some reason, looks almost like a quadratic equation - if you decide to substitute (cos(x))^2 with 1-(sin(x))^2

OpenStudy (anonymous):

you can then write sin(x) as x

OpenStudy (anonymous):

sin x=c u+2Kpi and (pi-u) +2Kpi cos x=c u+2kpi and -u +2Kpi tan x=c u+kpi here is the algorithm equation solutions

OpenStudy (anonymous):

1- x^2+5 x+5=0

OpenStudy (anonymous):

x=-1 or 6 x = Sin(theta)=-1 , or 6

OpenStudy (anonymous):

theta =270 + 2pi

OpenStudy (anonymous):

ok thats lookin easier now,

OpenStudy (anonymous):

thank you for your time imran